OpenStudy (anonymous):

if cosx=-5/13 and sinx<0, use half-angle identity to find cos(x/2)

OpenStudy (anonymous):

Since both cosine and sine are negative, the angle belongs to quadrant III.  Find the missing side. Then, use the identity, \[\cos^2\frac{x}{2}=\frac{1}{2}+\frac{1}{2}\cos x\]